A + 3.5-nC point charge is placed at one corner of a square (2.0 m on a side), and a - 5.7-nC charge is placed on a corner diagonally away from the first charge. What is the magnitude of the electric field at either of the two unoccupied corners?
Answer #1

field due to +3.5 nC is k*(3.5 nC)/(2.0 m)2 = 7.875 N/C
field due to -5.7 nC is k*(5.7 nC)/(2.0 m)2 = 12.825 N/C
So answer is √(7.8752 + 12.8252) = 15.05 N/C